OpenStudy (anonymous):

3.18x-4=8.72 Solve For x

OpenStudy (anonymous):

3.18x = 8.72 + 4 x = 12.72/3.18

OpenStudy (anonymous):

4

OpenStudy (anonymous):

carry out the division n u get the answer.....